OpenStudy (anonymous):

Can someone teach me how to factor a 5 termed polynomial? here is the polynomial function: 4x^4 + 5x^3 - 14x^2 - 4x +9

OpenStudy (amistre64):

if it is designed well enough, the rational roots thrm would apply to be able to knock it doen a degree of two

OpenStudy (amistre64):

if you solve the setup for "x" you can run some recursions to get approximations of a root.

OpenStudy (amistre64):

4x^4 + 5x^3 - 14x^2 - 4x +9 = 0 4x^4 + 5x^3 - 14x^2 +9 = 4x (4x^4 + 5x^3 - 14x^2 +9)/4 = x make a guess: g (4g^4 + 5g^3 - 14g^2 +9)/4 = x' then use x' as the new guess ... and it should start to converge to one of the roots

OpenStudy (anonymous):

I really don't get it. Can you explain it straight forward?

OpenStudy (anonymous):

and simple :) pretty please?

OpenStudy (amistre64):

the "simple" method relates to the rational roots thrm

OpenStudy (amistre64):

define the factors of the last term; and the first term coefficient divide the last by the first to develop a pool of possible roots to test out

OpenStudy (amistre64):

9 factors to: 1 3 9 4 factors to: 1 2 4 possible rational roots are therefore composed of some combination of: \[\pm\frac{1,3,9}{1,2,4}\]

OpenStudy (amistre64):

since the simplest option to test is "1" .... give it a shot when x=1, do we get a zero?

OpenStudy (anonymous):

there is no easy way to do this you have to find a zero first fortunately for you, you can check real fast that \(f(1)=0\) and then divide the polynomial by \(x-1\) to factor it

OpenStudy (amistre64):

by design, 1 is also a double root, which would simplify this down to the usual quadratic stuff

OpenStudy (amistre64):

... my eyeballing the graph may be in error tho ... looks lke a double lol

OpenStudy (anonymous):

i would use technology yeah, 1 is not a double root, just plays one on tv

OpenStudy (amistre64):

ugh, this thing was not "well designed" at all now was it

OpenStudy (anonymous):

no there is no good way to factor this in general there isn't so at least the question is honest

OpenStudy (anonymous):

yeah those were the numbers that sprang in to my head first thing

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

i was just checking by substitution that my guesses were right

OpenStudy (amistre64):

so, i spose this is a linear and a cubic prime?

OpenStudy (amistre64):

how can a cubic be prime? arent they spose to have at least one real root at all times ....

OpenStudy (anonymous):

\[(x-1) (4 x^3+9 x^2-5 x-9)\]

OpenStudy (anonymous):

this one has 4 roots usually "prime" means you can't factor using integers

OpenStudy (anonymous):

at least i think that is what "prime" means so for example \[x^2+4x-2\] is prime even though it has two zeros

OpenStudy (anonymous):

Hmm. I see, thank yooou for explaining. Now, i understand ^_^v Arigato gozaimas.

OpenStudy (anonymous):

Another one: how to factor this? x^3-7x+6 I factored it, i got (x^2-1) (x-6), but it's wrong.

OpenStudy (amistre64):

trial and error, try using x=1, or x=-1 as a simple test. there is also a opposite sign rule: on a continuous function; if f(a) is positive; and f(b) is negative ... then there has to be some value between a and b that produces a root

OpenStudy (anonymous):

Hmmm. I see, okaaay. Thanks again @amistre64 v^_^v